OpenStudy (xochital):

2x − y = 6 4x − 2y = 12 0 4 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system.

OpenStudy (anonymous):

I didn't get you what is it after 4x-2y=12

OpenStudy (xochital):

sorry, that's not very clear... What is the value of the x variable in the solution to the following system of equations? 2x − y = 6 4x − 2y = 12 0 4 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system.

OpenStudy (anonymous):

divide equation 2 by 2,u wil get same eqation

OpenStudy (anonymous):

no use in solving this 1

OpenStudy (xochital):

so no x value?

OpenStudy (anonymous):

for 2 variables,you need to "different" eqations to solve it

OpenStudy (anonymous):

right

OpenStudy (xochital):

thanks! :) can you help me with a few more?

OpenStudy (anonymous):

sure

OpenStudy (xochital):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1150. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many one-step equations eliminated today? 430 720 1150 1300

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

equations are easy

OpenStudy (anonymous):

1st is x+y=1150 and 2nd one is 3times x minus 2 times y that's 3x-2y=1300 now solve these equations

OpenStudy (anonymous):

got it?

OpenStudy (xochital):

hold on...

OpenStudy (badhi):

For the first question you've mentioned I have to make a correction.You have understood that there is one equation, ie. $$2x-y=6$$ take $$y=t \implies x=\frac{6+t}{2}$$ so, for each value t, there is infinite number of set of solutions. So, no solution for x is not the answer

OpenStudy (xochital):

oh okay, thank you :)

OpenStudy (anonymous):

by the way when you solve answer comes 720 cross check~

OpenStudy (anonymous):

:)

OpenStudy (xochital):

lol okay thanks, i just got it! :)

OpenStudy (anonymous):

good <3

OpenStudy (xochital):

how about... What is the value of the y variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11 −2 −3 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system

OpenStudy (anonymous):

multiply eq. 1 by 7 and 2 by 3, then subtract both,x will get eliminated

OpenStudy (anonymous):

y=-2 and x=-3 that's right,you got it?

OpenStudy (xochital):

yes! thank you again :D

OpenStudy (anonymous):

you are welcome :)

OpenStudy (anonymous):

any other?

OpenStudy (xochital):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 12 x – y = 3 0 −3 3 2

OpenStudy (anonymous):

multiply equation 2 by 2 and add both equations you will get x

OpenStudy (anonymous):

which is 3

OpenStudy (xochital):

lol okay i thought so. that's it, thank you so much :)

OpenStudy (anonymous):

no problem miss. :)